OpenStudy (anonymous):

How many solutions does the system of equations have? A. The system has no solution. B. The system has exactly one solution. C. The system has infinitely many solutions.

OpenStudy (anonymous):

i think it is a?

OpenStudy (anonymous):

@Muzzack

OpenStudy (anonymous):

By substitution method: To solve a pair of equations using substitution, first solve one of the equations for one of the variables. Then substitute the result for that variable in the other equation. (-1) y + 4x = 10 5y + 3x = 19 Choose one of the equations and solve it for y by isolating y on the left side of the equals sings. (-1) y + 4x = 10 Subtract 4x from both sides of the equation: (-1) y + 4x - 4x = 10 - 4x (-1) y = 10 - 4x Multiply both sides of the equation by -1: (-1) y * (-1) = (10 - 4x) (-1) y = -10 + 4x y = 4x - 10 Now substitute 4x - 10 for y in the other equation, 3x + 5y = 19: 5 (4x - 10) + 3x = 19 Multiply 5 times 4x - 10: 20x - 50 + 3x = 19 Add 20x to 3x (collect like terms): 23x - 50 = 19 Add 50 to both sides of the equation: 23x - 50 + 50 = 19 + 50 23x = 69 Divide both sides of the equation by 23: 23x / 23 = 69 / 23 x = 69 / 23 x = 3 Now substitute 3 for x in y = 4x - 10. Since the resulting equation contains only one variable, you can solve for "y" directly: y = 4 * 3 - 10 y = 12 - 10 y = 2 Solution: x = 3 y = 2 It just have one answer

OpenStudy (anonymous):

so its b?

OpenStudy (anonymous):

@wakeboarder13

OpenStudy (anonymous):

so its b? @ikatouni

OpenStudy (anonymous):

@ikatouni is it B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes @wakeboarder13